Date:
Tuesday, September 8, 2015
Telescope:
CTIO Blanco 4-m
Submitter:
Ross Cawthon
Submitter Email:
Observer Support:
Manuel Hernandez
Night Assistant:
Claudio Aguilera
Also Present:
James Lasker, David James, Camila Naverrete
Problems:
Clouds.
Program:
Proposal:
2012B-0001
Target of Oportunity:
No
Time Observed:
1.5
Time Engineering:
0.0
Time lost technical:
0.0
Time lost weather:
3.5
Time lost other:
0.0
Description:
Dark Energy Survey
Primary Investigator:
Josh Frieman
Email:
Institution:
Fermilab
Instrument:
CTIO 4m DECam
Organization:
NOAO
Total Program Time:
5
Proposal:
2015B-0601
Target of Oportunity:
No
Time Observed:
0.0
Time Engineering:
0.0
Time lost technical:
0.0
Time lost weather:
5.0
Time lost other:
0.0
Description:
DECam Search for the Elusive Stellar Counterpart of the Magellanic Stream
Primary Investigator:
J. Chaname
Email:
Institution:
Pontificia Universidad Catolica de Chile
Instrument:
CTIO 4m DECam
Organization:
Chile
Total Program Time:
5
Total Times | Time Observed | Time Engineering | Time lost technical | Time lost weather | Time lost other | Total Program Time |
---|---|---|---|---|---|---|
2015-09-08 | 1.5 | 0.0 | 0.0 | 8.5 | 0.0 | 10 |
*Observing Team*
_DES Observers:_ Ross Cawthon, James Lasker, David James
_Telops:_ Manuel Hernandez, Claudio Aguilera
_Remote Support:_ ---
*Observing Plan*
# standard bias and dome flat sequence
# evening standard stars
# run ObsTac queue
# Handoff to Community Observers at 23:30 CLT
# Take over from Community Observers at 04:30 CLT
# end of night calibrations
*Conditions*
# weather in afternoon: 4pm (local): Clear but with forecast of partial clouds
# weather after dinner: 6pm (local): Partly Cloudy
# weather after sunset: 8pm (local): Clouds covering most of sky to varying degrees of opacity. Small circle overhead without clouds. Standards taken but likely low quality.
# weather during night: 9:15pm (local): Rasicam shows clouds encroaching from the East. Significant portions of DES Fields under some amount of cloud cover.
# Rest of the night under significant cloud cover, stars detected dropped to zero. Closed dome at 11:30 CLT and did not reopen. Zero patches of sky clear.
*Accomplishments*
# Afternoon: zeros done, (ugrizYVR) dome flats done, checked, OK
# Evening: Standards (E8-A, E6-A, High Airmass Standards skipped due to total cloud cover)
# aTmCam and GPS Monitor Checked.
# First Half Survey Exposures 472562-472645 (mostly r and i, few g and z scattered throughout, 6 Y at start of night, all likely spoiled by clouds);
# teff of -1.00 (Exposures 472604-5)
*Problems tagged as Synopsis_Problem*
# None
*Minor Problems and Relevant Information*
# Issue on DECals, see night summary.
*Notes for Data Management*
# Due to Clouds, it might be worth declaring exposures 472552-472561(Standards) as bad
# Due to Clouds causing teff near or equal to 0.00 for 472563-472645 (Survey Exposures), might be worth marking them as bad.
# Due to guide star failures and/or quickreduce failures, exposures 472609-614 and 472617-645 are almost certainly bad
# DECal issue